In a video game, a specific "attack" succeeds at a rate of 18% and every try is independent. Detemine the probability that the first "attack" lands after the 4th try.
Answer #1

18% success sate for Specific affack an 1-P=0.82 p=0.18 every toy is Inderendent plista attack tands after the ith try) a BobI method P-0-18 1--0.62 Plafter with toy) = p(xsu) PvAbiy P(1st affack on ether yor56.-) (x25)+(K-() exey 4 ((-p) p tipfe th